# Restore Text First, Enhance Image Later: Two-Stage Scene Text Image Super-Resolution with Glyph Structure Guidance

### v1

@@ 3.3 Image Enhancement stage
Real-ESRGAN 22.30 0.787 0.239 0.188 53.60 23.99 0.790 0.248 0.194 30.60
HAT 23.61 0.830 0.214 0.176 51.16 25.17 0.815 0.249 0.198 30.12
MARCONet 21.89 0.785 0.238 0.150 52.97 22.13 0.768 0.306 0.205 34.28
SeeSR 23.59 0.822 0.195 0.169 43.75 23.64 0.788 0.219 0.184 26.73
SupIR 21.78 0.723 0.310 0.198 44.94 23.62 0.754 0.308 0.207 27.91
DiffTSR 22.10 0.768 0.278 0.168 44.91 22.41 0.767 0.300 0.189 31.25
DiffBIR 22.44 0.747 0.260 0.201 46.44 23.67 0.724 0.262 0.197 23.10
OSEDiff 21.86 0.771 0.197 0.127 41.00 25.07 0.819 0.201 0.169 20.53
DreamClear 22.47 0.772 0.216 0.157 38.97 24.10 0.773 0.238 0.191 21.75
TSD-SR 21.43 0.754 0.220 0.175 47.21 22.79 0.757 0.207 0.194 24.08
DiT4SR 20.54 0.764 0.268 0.186 49.79 23.16 0.767 0.215 0.159 20.58
TADiSR 23.83 0.790 0.286 0.154 44.42 24.61 0.796 0.203 0.160 36.61
@@
Table 3: Evaluation of image quality and text accuracy. Metrics with cr{cr} are computed only on text regions. TIGER (ours) achieves the best performance in both.
Methods Real-CE UltraZoom-ST
PSNRcr SSIMcr LPIPScr DISTScr OCR-A PSNRcr SSIMcr LPIPScr DISTScr OCR-A
Real-ESRGAN 21.71 0.824 0.257 0.224 0.560 21.25 0.786 0.302 0.266 0.371
HAT 23.03 0.862 0.249 0.239 0.566 22.18 0.813 0.291 0.276 0.379
MARCONet 20.68 0.786 0.267 0.223 0.550 17.48 0.690 0.529 0.366 0.334
SeeSR 22.74 0.844 0.237 0.212 0.374 20.45 0.767 0.297 0.241 0.191
SupIR 19.96 0.778 0.326 0.263 0.278 20.10 0.756 0.344 0.287 0.236
DiffTSR 20.46 0.818 0.275 0.239 0.441 17.00 0.665 0.452 0.313 0.317
DiffBIR 21.20 0.792 0.278 0.224 0.376 20.70 0.757 0.302 0.245 0.266
OSEDiff 19.38 0.768 0.269 0.208 0.268 21.43 0.798 0.276 0.263 0.300
DreamClear 22.74 0.845 0.190 0.168 0.502 20.50 0.768 0.317 0.276 0.214
TSD-SR 19.33 0.763 0.279 0.234 0.349 19.42 0.748 0.296 0.249 0.266
DiT4SR 17.95 0.738 0.317 0.244 0.292 19.73 0.760 0.273 0.216 0.237
TADiSR 23.39 0.855 0.253 0.258 0.647 21.59 0.799 0.360 0.336 0.384
@@
4 Dataset and Benchmark
Existing datasets like Ma et al. (2023) offer only subtle degradation, as shown in Fig. 3 (a), making them insufficient for evaluating model robustness. To address the lack of both challenging and publicly available datasets tailored for scene text image super-resolution, especially for Chinese text, we introduce UltraZoom-ST, a challenging real-world benchmark. It is collected using ViVO X200 Ultra equipped with four fixed focal lengths (14 mm, 35 mm, 85 mm, 200 mm), enabling image pairs with extreme $\times 14.29$ zoom. It includes diverse scenes—street views, book covers, advertisements, menus, and posters—captured under varied lighting, with all text lines manually annotated. However, such extreme zoom introduces severe misalignment that breaks pixel-wise optimization methods like Cai et al. (2019), originally designed for moderate $\times 2$ to $\times 4$ zoom. To overcome this, we design a Cascade Coarse-to-fine Alignment pipeline shown in Fig. 3 (b): images are first sorted by focal length, and each low-quality image is sequentially aligned to its next higher-focal neighbor using either optimization-based registration or feature-based extractors (e.g., Lowe (1999); Künzel et al. (2025)), then refined against the 200 mm ground truth for precise alignment.
@@ 4 Dataset and Benchmark
Through meticulous annotating and aligning, we obtained a total of 5,036 image pairs, with 49,675 lines of text. We set images under 200mm focal lengths as GT, and obtain 1,439, 1,798, and 1,799 pairs for $\times 14.29$, $\times 5.71$, and $\times 2.35$ zooming modes, respectively. Among them, we randomly select 470, 589, and 581 pairs for evaluation under each zooming mode. Each image pair contains one or more text lines. These evaluation sets enable us to evaluate the performance of models under more complex and challenging scenarios.
@@
For the UltraZoom-ST-benchmark, we utilize 5 evaluation metrics to examine the quality of image super-resolution on the scale of the full image. Firstly, we adopt Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) (Wang et al., 2019), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), Deep Image Structure and Texture Similarity (DISTS) (Ding et al., 2020), and Fréchet Inception Distance (FID) (Heusel et al., 2017) for evaluation of image quality. Specifically, PSNR and SSIM are computed in the pixel space to quantify low-level reconstruction fidelity, LPIPS and DISTS are computed in the feature space to assess perceptual similarity, and FID is employed to evaluate the distributional discrepancy between generated and real images. However, these metrics do not directly reflect the quality or accuracy of super-resolved text. To address this, we crop annotated text regions from the images and compute PSNR, SSIM, LPIPS, and DISTS on these cropped regions, denoted as $\text{PSNR}{cr}$, $\text{SSIM}{cr}$, $\text{LPIPS}{cr}$, and $\text{DISTS}{cr}$. For text accuracy, we apply an OCR model to recognize the text and compare its outputs with the ground-truth annotations using the Levenshtein ratio (Yujian & Bo, 2007):
$\text{OCR-A}=(\text{Len}(s_{pred})+\text{Len}(s_{gt})-\text{Dist}(s_{pred},s_{gt}))/({\text{Len}(s_{pred})+\text{Len}(s_{gt})}),$ (8)
@@
where $s_{pred}$ stands for the predicted text sequence, $s_{gt}$ denotes the annotated ground-truth text sequence, and $\text{Dist}(\cdot,\cdot)$ is the Levenshtein distance between the text sequences.
Figure 4: Qualitative Evaluation on Real-CE and UltraZoom-ST.
@@ 5.1 Implementation Details
We combine synthetic data (built upon LSDIR (Li et al., 2023b) with text rendered via LBTS (Tang et al., 2023) and Real-ESRGAN (Wang et al., 2021) degradation) and real paired data from Real-CE (Ma et al., 2023) and UltraZoom-ST. Misaligned pairs in Real-CE are filtered and reannotated, yielding 337 training and 188 testing pairs. Stage 1 uses cropped text regions and is trained using an IDM-based architecture (Zhang et al., 2024a) with combined segmentation and reconstruction losses. Stage 2 is based on Stable Diffusion 3.5 (Esser et al., 2024) and employs a tile-based inference strategy (Yu et al., 2024; Hu et al., 2025); it is pretrained on synthetic data and fine-tuned on real data. More detailed information can be found in the Appendix A.3.
@@ 5.2.1 Quantitative Results
We evaluate existing competing methods, including GAN-based image super-resolution approaches such as Real-ESRGAN (Wang et al., 2021) and HAT (Chen et al., 2023); diffusion-based approaches such as SeeSR (Wu et al., 2024b), SupIR (Yu et al., 2024), DiffBIR (Lin et al., 2024), OSEDiff (Wu et al., 2024a), DreamClear (Ai et al., 2024), TSD-SR (Dong et al., 2025b), and DiT4SR (Duan et al., 2025); as well as text-focused reconstruction methods such as MARCONet (Li et al., 2023a), DiffTSR (Zhang et al., 2024a), and TADiSR (Hu et al., 2025). Evaluations are performed on the Real-CE Benchmark (Ma et al., 2023) and the benchmark described in Sec. 4. To ensure fairness, we fine-tune the released pre-trained models on the training sets of both benchmarks using the official code when available. Following Hu et al. (2025), we integrate the outputs of MARCONet and DiffTSR with HAT-generated results to simulate real-world application scenarios and enable comprehensive full-image evaluation. We evaluate Real-CE at the hardest difficulty level ($\times 4$), while UltraZoom-ST is evaluated across all difficulty levels—with average results reported here. Detailed per-level evaluations appear in the Appendix A.4.
As shown in Table 2 and Table 3 , TIGER outperforms competing methods on both the Real-CE and UltraZoom-ST benchmarks in terms of image quality and text accuracy. TADiSR, limited by the resolution of its cross-attention mechanism, performs poorly under severe degradation and with small text. DiffTSR and MARCONet fail to effectively handle text backgrounds and struggle with text regions that have a large width-to-height aspect ratio. In contrast, our method restores images with high quality and fine-grained text, owing to the decoupling strategy. It achieves OCR-A scores above 0.67 and 0.43, and demonstrates overall superiority in both pixel-level and perceptual accuracy.
@@ 5.2.2 Qualitative Results
Fig. 4 shows visual comparisons on Real-CE and UltraZoom-ST. Real-world degradations cause noise, blurred strokes, and distortions in the input images. GAN-based methods reduce noise but fail to fix text structure, while diffusion-based SR methods, though powerful, lack text-structure guidance and often distort strokes, sometimes making text unreadable.
@@ 5.3 Ablation Study
Table 4: Validation on UltraZoom-ST (UZST). Finetuning improves OCR-A, proving effectiveness.
@@
OSEDiff w/o UZST 23.50 0.791 0.197 0.162 24.30 0.228
OSEDiff w/   UZST 25.07 0.819 0.201 0.169 20.53 0.300
DiT4SR w/o UZST 22.58 0.754 0.252 0.196 26.93 0.193
DiT4SR w/   UZST 23.17 0.767 0.215 0.159 20.58 0.237
Ours w/o UZST 22.96 0.782 0.341 0.253 33.78 0.400
Ours w/   UZST 25.48 0.830 0.196 0.156 20.01 0.430
@@
Std Font Guidance 23.31 0.786 0.249 0.164 41.83 0.553
SAM-TS Extraction 23.91 0.835 0.212 0.147 41.22 0.579
LDM Guidance 22.42 0.798 0.222 0.147 43.08 0.601
@@
Effectiveness of the training set. Table 4 reports consistent OCR-A improvements across different architectures, including the UNet-based OSEDiff and the DiT-based DiT4SR. This demonstrates that even models without explicit mechanisms for text structure modeling can still learn the mapping from low- to high-resolution text using our dataset, further validating its effectiveness. Due to their lack of a dedicated mechanism for text structure modeling, they cannot outperform our methods.
Effectiveness of the TIGER component. We set stage 2 as the fixed baseline for image restoration and validate the effectiveness of our framework through 3 ablated variants:
  • Standard Font Guidance. As shown in Fig. 5 (first column), the standard font, while structurally correct, provides weak guidance due to style and position mismatches. As a result, the restored text exhibits poor accuracy compared to TIGER, with the OCR-A dropping from 0.671 to 0.553, as shown in Table 5.
  • SAM-TS Extraction Hu et al. (2025); Ye et al. (2024). While this improves visual quality by better aligning output masks with the LR image, SAM-TS can only extract distorted structures and can’t recover degraded text (Fig. 5, second column), leading to suboptimal performance compared with TIGER, with OCR-A dropping from 0.671 to 0.579.
  • LDM Guidance. This variant uses a latent diffusion model conditioned on LR to restore text structure, achieving better accuracy by reconstructing or compensating for lost text. However, without separate learning from synthetic and real-world data, it struggles with high-quality text structures in real-world scenarios. The masks remain noisy under real-world degradations (Fig. 5, third column), leading to poorer accuracy compared to TIGER, with OCR-A dropping from 0.671 to 0.601 as shown in Table 5.
Our method leverages RGB output to jointly learn from real-world degradations and accurate synthetic masks, producing high-quality masks and achieving the highest accuracy overall.
@@ 6 Conclusion
In this paper, we delve into the extensively researched problem of scene text super-resolution. To address this challenge, we propose a novel approach called TIGER, which decouples the glyph structure restoration and image enhancement. For the restoration of text structure, we propose the text restoration pipeline, which enables a 2-phase training strategy to fully take advantage of both synthetic and real-world data. For the image enhancement, we propose the image enhancement pipeline, which effectively utilizes the glyph structure to restore both text fidelity and image details coherently in full-image super-resolution. In terms of training data and benchmark, we present the UltraLens-ST, the first real-world scene text dataset that contains extreme zooming mode, offering more challenging scenarios for the field. Extensive experiments on Real-CE and UltraLens-ST demonstrate the superiority of TIGER over existing methods.
Reproducibility Statement
To ensure reproducibility, we have made the following efforts: (1) We will release our code and dataset. (2) We provide implementation details in Sec. 5.1 and Appendix A.3, including the training process and selection of hyper-parameters. (3) We provide details on evaluation metrics and dataset preparation in Sec. 4 and Appendix A.2, and the code and data will be made available along with it.
Ethics Statement
This work focuses on improving scene text image super-resolution to support beneficial applications such as enhancing accessibility, document restoration, and navigation assistance. However, we acknowledge potential risks, including misuse in privacy-sensitive contexts (e.g., recovering text from personal or social media images) or unintended deployment in surveillance. To mitigate such risks, users are encouraged to combine our methods with privacy-preserving techniques such as watermarking or selective inpainting. Our UltraZoom-ST dataset was collected from public, non-sensitive scenes under ethical guidelines, without including personally identifiable or private information. We believe the societal benefits of improved text image restoration outweigh potential risks, provided that the technology is applied responsibly.
Appendix A Appendix
@@
85mm ($\times 2.35$)
Real-ESRGAN 25.84 0.854 0.139 0.130 28.25 23.59 0.843 0.170 0.168 0.632
HAT 27.33 0.873 0.128 0.133 27.21 24.76 0.870 0.139 0.146 0.643
MACRONet 22.90 0.804 0.234 0.173 40.19 17.92 0.693 0.515 0.360 0.542
SeeSR 25.02 0.826 0.167 0.159 32.31 21.62 0.796 0.245 0.207 0.374
SupIR 25.58 0.823 0.202 0.160 29.33 21.91 0.808 0.239 0.214 0.351
DiffTSR 23.48 0.805 0.230 0.157 35.77 17.92 0.688 0.392 0.285 0.482
DiffBIR 25.25 0.793 0.175 0.152 24.57 22.26 0.803 0.218 0.190 0.391
OSEDiff 26.38 0.852 0.153 0.137 23.92 22.84 0.833 0.200 0.204 0.433
DreamClear 25.31 0.811 0.169 0.151 24.19 21.44 0.790 0.253 0.221 0.291
TSD-SR 23.99 0.800 0.161 0.165 30.40 20.62 0.786 0.232 0.218 0.401
DiT4SR 24.71 0.810 0.173 0.143 29.15 21.25 0.800 0.221 0.193 0.482
TADiSR 26.31 0.858 0.177 0.149 30.47 23.55 0.856 0.175 0.178 0.581
Ours 27.00 0.871 0.123 0.120 22.34 24.31 0.860 0.135 0.142 0.658
35mm ($\times 5.71$)
Real-ESRGAN 23.70 0.785 0.218 0.184 42.62 20.86 0.782 0.272 0.235 0.348
HAT 25.05 0.819 0.207 0.182 41.32 21.98 0.816 0.250 0.241 0.354
MACRONet 22.28 0.774 0.284 0.199 49.14 17.43 0.694 0.523 0.361 0.316
SeeSR 23.76 0.795 0.198 0.175 37.92 20.55 0.778 0.271 0.222 0.317
SupIR 23.49 0.746 0.302 0.204 41.80 20.00 0.757 0.335 0.282 0.248
DiffTSR 22.57 0.773 0.281 0.184 45.78 16.99 0.671 0.443 0.312 0.312
DiffBIR 23.76 0.731 0.245 0.196 35.99 20.82 0.770 0.275 0.234 0.293
OSEDiff 25.35 0.827 0.192 0.166 29.06 21.62 0.809 0.263 0.263 0.307
DreamClear 24.20 0.778 0.209 0.177 32.08 20.66 0.779 0.283 0.251 0.247
TSD-SR 22.93 0.757 0.200 0.199 37.61 19.64 0.760 0.282 0.246 0.281
DiT4SR 23.45 0.774 0.204 0.158 31.73 20.12 0.779 0.250 0.208 0.266
TADiSR 24.68 0.795 0.362 0.227 52.13 21.57 0.798 0.366 0.347 0.375
Ours 25.63 0.834 0.185 0.150 28.82 22.14 0.823 0.204 0.194 0.456
14mm ($\times 14.29$)
Real-ESRGAN 22.09 0.718 0.423 0.284 89.22 18.87 0.721 0.503 0.427 0.116
HAT 22.66 0.738 0.451 0.299 89.58 19.27 0.738 0.533 0.484 0.123
MACRONet 21.00 0.716 0.422 0.254 79.67 17.01 0.680 0.553 0.382 0.123
SeeSR 21.79 0.731 0.310 0.228 64.31 18.87 0.717 0.396 0.308 0.157
SupIR 21.38 0.681 0.447 0.268 70.09 17.99 0.690 0.487 0.383 0.110
DiffTSR 20.88 0.713 0.412 0.233 74.07 15.87 0.630 0.538 0.348 0.120
DiffBIR 21.61 0.630 0.392 0.255 69.64 18.61 0.685 0.438 0.328 0.114
OSEDiff 23.11 0.768 0.274 0.214 49.76 19.44 0.740 0.385 0.334 0.127
DreamClear 22.45 0.719 0.364 0.260 62.95 19.14 0.725 0.440 0.377 0.124
TSD-SR 21.14 0.705 0.273 0.225 57.07 17.67 0.686 0.393 0.293 0.119
DiT4SR 20.88 0.707 0.282 0.181 47.45 17.36 0.687 0.367 0.253 0.117
TADiSR 22.43 0.721 0.555 0.320 98.44 19.20 0.730 0.585 0.518 0.152
Ours 23.41 0.774 0.301 0.209 54.74 19.73 0.748 0.374 0.322 0.160
Total
RealEsrGAN 23.99 0.790 0.248 0.194 30.60 21.25 0.786 0.302 0.266 0.371
HAT 25.17 0.815 0.249 0.198 30.12 22.18 0.813 0.291 0.276 0.379
MARCONet 22.13 0.768 0.306 0.205 34.28 17.48 0.690 0.529 0.366 0.334
SeeSR 23.64 0.788 0.219 0.184 26.73 20.45 0.767 0.297 0.241 0.191
SupIR 23.62 0.754 0.308 0.207 27.91 20.10 0.756 0.344 0.287 0.236
DiffTSR 22.41 0.767 0.300 0.189 31.25 17.00 0.665 0.452 0.313 0.317
DiffBIR 23.67 0.724 0.262 0.197 23.10 20.70 0.757 0.302 0.245 0.266
OSEDiff 25.07 0.819 0.201 0.169 20.53 21.43 0.798 0.276 0.263 0.300
DreamClear 24.10 0.773 0.238 0.191 21.75 20.50 0.768 0.317 0.276 0.214
TSD-SR 22.79 0.757 0.207 0.194 24.08 19.42 0.748 0.296 0.249 0.266
DiT4SR 23.16 0.767 0.215 0.159 20.58 19.73 0.760 0.273 0.216 0.237
TADiSR 24.61 0.796 0.203 0.160 36.61 21.59 0.799 0.360 0.336 0.384
Ours 25.48 0.830 0.196 0.156 20.01 22.22 0.814 0.228 0.212 0.430
@@
HAT 6670.32 1086.68 0.379
DiffTSR 58502.14 8610.59 0.317
DiT4SR w/o llava 160787.14 17385.14 0.237
DreamClear w/o llava 412843.67 83193.81 0.214
TADiSR 4497.96 342.32 0.384
TIGER (ours) Stage 1 6215.85 663.5 0.430

### v2

@@ 4 Dataset and Benchmark
Real-ESRGAN 22.30 0.787 0.239 0.188 53.60 56.0% 23.99 0.790 0.248 0.194 30.60 37.1%
HAT 23.61 0.830 0.214 0.176 51.16 56.6% 25.17 0.815 0.249 0.198 30.12 37.9%
MARCONet 21.89 0.785 0.238 0.150 52.97 55.0% 22.13 0.768 0.306 0.205 34.28 33.4%
SeeSR 23.59 0.822 0.195 0.169 43.75 37.4% 23.64 0.788 0.219 0.184 26.73 22.8%
SupIR 21.78 0.723 0.310 0.198 44.94 27.8% 23.62 0.754 0.308 0.207 27.91 23.6%
DiffTSR 22.10 0.768 0.278 0.168 44.91 44.1% 22.41 0.767 0.300 0.189 31.25 31.7%
DiffBIR 22.44 0.747 0.260 0.201 46.44 37.6% 23.67 0.724 0.262 0.197 23.10 26.6%
OSEDiff 21.86 0.771 0.197 0.127 41.00 26.8% 25.07 0.819 0.201 0.169 20.53 28.9%
DreamClear 22.47 0.772 0.216 0.157 38.97 50.2% 24.10 0.773 0.238 0.191 21.75 21.4%
TSD-SR 21.43 0.754 0.220 0.175 47.21 34.9% 22.79 0.757 0.207 0.194 24.08 26.6%
DiT4SR 20.54 0.764 0.268 0.186 49.79 29.2% 23.16 0.767 0.215 0.159 20.58 23.7%
TADiSR 23.83 0.790 0.286 0.154 44.42 64.7% 24.61 0.796 0.203 0.160 36.61 36.6%
@@
Table 3: Evaluation of image quality on text regions and text accuracy compared to LR ($\Delta$ OCR-A). TIGER achieves the best performance.
Methods Real-CE UZ-ST
@@
Real-ESRGAN 21.71 0.824 0.257 0.224 -8.8% 21.25 0.786 0.302 0.266 -4.7%
HAT 23.03 0.862 0.249 0.239 -8.2% 22.18 0.813 0.291 0.276 -3.9%
MARCONet 20.68 0.786 0.267 0.223 -9.8% 17.48 0.690 0.529 0.366 -8.3%
SeeSR 22.74 0.844 0.237 0.212 -27.4% 20.45 0.767 0.297 0.241 -18.9%
SupIR 19.96 0.778 0.326 0.263 -37.0% 20.10 0.756 0.344 0.287 -18.2%
DiffTSR 20.46 0.818 0.275 0.239 -20.7% 17.00 0.665 0.452 0.313 -10.0%
DiffBIR 21.20 0.792 0.278 0.224 -27.2% 20.70 0.757 0.302 0.245 -15.2%
OSEDiff 19.38 0.768 0.269 0.208 -38.0% 21.43 0.798 0.276 0.263 -12.8%
DreamClear 22.74 0.845 0.190 0.168 -14.6% 20.50 0.768 0.317 0.276 -20.3%
TSD-SR 19.33 0.763 0.279 0.234 -29.9% 19.42 0.748 0.296 0.249 -15.1%
DiT4SR 17.95 0.738 0.317 0.244 -35.6% 19.73 0.760 0.273 0.216 -18.0%
TADiSR 23.39 0.855 0.253 0.258 -1.0% 21.59 0.799 0.360 0.336 -5.1%
[…]
@@
Through meticulous annotation and alignment, we obtained 5,036 image pairs with 49,675 text lines. We treat images captured under 200 mm focal lengths as GT, yielding 1,439, 1,798, and 1,799 pairs for the $\times 14.29$, $\times 5.71$, and $\times 2.35$ zoom modes, respectively. Among them, we randomly select 470, 589, and 581 pairs for evaluation under each zoom mode. Each image pair contains one or more text lines. These evaluation sets allow us to assess model performance in more complex and challenging scenarios.
@@
For the UZ-ST-benchmark, we utilize 5 evaluation metrics to examine the quality of image super-resolution on the scale of the full image. Firstly, we adopt Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [wang2019textsr], Learned Perceptual Image Patch Similarity (LPIPS) [zhang2018unreasonable], Deep Image Structure and Texture Similarity (DISTS) [ding2020image], and Fréchet Inception Distance (FID) [heusel2017gans] for evaluation of image quality. Specifically, PSNR and SSIM are computed in the pixel space to quantify low-level reconstruction fidelity, LPIPS and DISTS are computed in the feature space to assess perceptual similarity, and FID is employed to evaluate the distributional discrepancy between generated and real images. However, these metrics do not directly reflect the quality or accuracy of super-resolved text. To address this, we crop annotated text regions from the images and compute PSNR, SSIM, LPIPS, and DISTS on these cropped regions, denoted as $\text{PSNR}{cr}$, $\text{SSIM}{cr}$, $\text{LPIPS}{cr}$, and $\text{DISTS}{cr}$. For text accuracy, we apply an OCR model to recognize the text and compare its outputs with the ground-truth annotations using the Levenshtein ratio [yujian2007normalized]:
$\text{OCR-A}=\frac{\text{Len}(s_{pred})+\text{Len}(s_{gt})-\text{Dist}(s_{pred},s_{gt})}{\text{Len}(s_{pred})+\text{Len}(s_{gt})},$ (8)
@@
where $s_{pred}$ stands for the predicted text sequence, $s_{gt}$ denotes annotated ground-truth text sequence, and $\text{Dist}(\cdot,\cdot)$ is the Levenshtein distance between text sequences.
Figure 4: Qualitative Evaluation on Real-CE and UZ-ST.
@@ 5.1 Implementation Details
We combine synthetic data (built upon LSDIR [li2023lsdir] with text rendered via LBTS [tang2023scene] and Real-ESRGAN [wang2021real] degradation) and real data from Real-CE [ma2023benchmark] and UZ-ST. Misaligned pairs in Real-CE are filtered and reannotated, yielding 337 training and 188 testing pairs. Stage 1 uses cropped text regions and is trained using an IDM-based architecture [zhang2024diffusion] with combined segmentation and reconstruction losses. Stage 2 is based on Stable Diffusion 3.5 [esser2024scaling] and employs a tile-based inference strategy [yu2024scaling, hu2025text]; it is pretrained on synthetic data and fine-tuned on real data. For more details, see the Sec. 11.
@@ 5.2.1 Quantitative Results
We evaluate existing competing methods, including GAN-based image super-resolution approaches such as Real-ESRGAN [wang2021real] and HAT [chen2023activating]; diffusion-based approaches such as SeeSR [wu2024seesr], SupIR [yu2024scaling], DiffBIR [lin2024diffbir], OSEDiff [wu2025one], DreamClear [ai2024dreamclear], TSD-SR [dong2025tsd], and DiT4SR [duan2025dit4sr]; as well as text-focused reconstruction methods such as MARCONet [li2023learning], DiffTSR [zhang2024diffusion], and TADiSR [hu2025text]. Evaluations are performed on the Real-CE Benchmark [ma2023benchmark] and the benchmark described in Sec. 4. To ensure fairness, we fine-tune the released pre-trained models on the training sets of both benchmarks using the official code when available. Following [hu2025text], we integrate the outputs of MARCONet and DiffTSR with HAT-generated results to simulate real-world application scenarios and enable comprehensive full-image evaluation. We evaluate Real-CE at the hardest difficulty level ($\times 4$), while UZ-ST is evaluated across all difficulty levels—with average results reported here. Detailed per-level evaluations appear in the Sec. 14.
As shown in Tab. 2 and Tab. 3 , TIGER outperforms competing methods on both the Real-CE and UZ-ST benchmarks in terms of image quality and text accuracy. TADiSR, limited by the resolution of its cross-attention mechanism, performs poorly under severe degradation and with small text. DiffTSR and MARCONet fail to effectively handle text backgrounds and struggle with text regions that have a large width-to-height aspect ratio. In contrast, our method restores images with high quality and fine-grained text, owing to the decoupling strategy. It achieves OCR-A scores above 0.67 and 0.43, and demonstrates overall superiority in both pixel-level and perceptual accuracy.
@@ 5.2.2 Qualitative Results
Fig. 4 shows visual comparisons on Real-CE and UZ-ST. Real-world degradations cause noise, blurred strokes, and distortions in the input images. GAN-based methods reduce noise but fail to fix text structure, while diffusion-based SR methods, though powerful, lack text-structure guidance and often distort strokes, sometimes making text unreadable.
@@ 5.3 Ablation Study
Table 4: Validation on the effectiveness of UZ-ST. Finetuning improves OCR-A, proving effectiveness.
@@
OSEDiff w/o UZ-ST 23.50 0.791 0.197 0.162 24.30 22.8%
OSEDiff w/   UZ-ST 25.07 0.819 0.201 0.169 20.53 28.9%
DiT4SR w/o UZ-ST 22.58 0.754 0.252 0.196 26.93 19.3%
DiT4SR w/   UZ-ST 23.17 0.767 0.215 0.159 20.58 23.7%
Ours w/o UZ-ST 22.96 0.782 0.341 0.253 33.78 40.0%
Ours w/   UZST 25.48 0.830 0.196 0.156 20.01 43.0%
Table 5: Ablation on OCR prediction. The model remains strong with null or random OCR text, indicating limited reliance on OCR.
[…]
TADiSR 24.68 0.795 0.362 0.227 52.13 35.5%
Null Text 25.59 0.834 0.188 0.151 28.93 40.4%
Random Text 25.60 0.833 0.188 0.150 28.92 40.3%
Predicted Text (Ours) 25.63 0.834 0.185 0.150 28.82 44.6%
Effectiveness of the training set. Tab. 4 reports consistent OCR-A improvements across different architectures, including the UNet-based OSEDiff and the DiT-based DiT4SR. This demonstrates that even models without explicit mechanisms for text structure modeling can still learn the mapping from low- to high-resolution text using our dataset, further validating its effectiveness. Due to their lack of a dedicated mechanism for text structure modeling, they cannot outperform our methods.
OCR Influence. To investigate how OCR prediction affects the accuracy of our framework, we set the predicted text by OCR in stage 1 as null text or random text and test them on the 35mm subset of UZ-ST. As shown in Tab. 5, even without the OCR model’s predicted text, our method attains a strong OCR-A score of 40.4%, outperforming TAiDSR. This demonstrates that our model in stage 1 benefits from the inherent text structure in the LR input to restore text structure and is not overly dependent on the OCR model.
Effectiveness of the TIGER component. We set stage 2 as the fixed baseline for image restoration and validate the effectiveness of our framework through 4 ablated variants:
  • Empty Mask Guidance. As shown in Tab. 4, empty masks provide zero structure guidance, and the model can only rely on the LR for text restoration. Therefore, the OCR-A drops from 67.3% to 59.5%.
  • Standard Font Guidance. As shown in the second column of Fig. 5, the standard font, while structurally correct, provides weak guidance due to style and position mismatches. As a result, the restored text exhibits poor accuracy compared to TIGER, with the OCR-A dropping from 67.1% to 55.3%, as shown in Tab. 6.
  • SAM-TS Extraction [hu2025text, ye2024hi]. While this improves visual quality by better aligning output masks with the LR image, SAM-TS can only extract distorted structures and can’t recover degraded text (Fig. 5, third column), leading to suboptimal performance compared with TIGER, with OCR-A dropping from 0.671 to 0.579.
  • LDM Guidance. This variant uses a latent diffusion model conditioned on LR to restore text structure, achieving better accuracy by reconstructing or compensating for lost text. However, without separate learning from synthetic and real-world data, it struggles with high-quality text structures in real-world scenarios. As shown in Fig. 5, fourth column, the masks remain noisy under real-world degradations, leading to lower accuracy than TIGER, with OCR-A dropping from 67.1% to 60.1% as reported in Tab. 6.
Our method uses RGB output to jointly learn from real-world degradations and accurate synthetic masks, yielding high-quality masks and achieving the highest accuracy.
@@
Empty Mask 23.33 0.792 0.225 0.172 42.93 59.5%
Std Font Guidance 23.31 0.786 0.249 0.164 41.83 55.3%
SAM-TS Extraction 23.91 0.835 0.212 0.147 41.22 57.9%
LDM Guidance 22.42 0.798 0.222 0.147 43.08 60.1%
@@ 6 Conclusion
In this paper, we delve into the extensively researched problem of scene text super-resolution. To address this challenge, we propose a novel approach, TIGER, that decouples glyph structure restoration from image enhancement. For the restoration of text structure, we propose the text restoration pipeline, which enables a 2-phase training strategy to fully take advantage of both synthetic and real-world data. For the image enhancement, we propose the image enhancement pipeline, which effectively utilizes the glyph structure to restore both text fidelity and image details coherently in full-image super-resolution. In terms of training data and benchmark, we present the UltraZoom-ST, the first real-world scene text dataset that contains extreme zooming mode, offering more challenging scenarios for the field. Extensive experiments on Real-CE and UZ-ST demonstrate the superiority of TIGER over existing methods.
@@
Supplementary Material
@@
Table 8: Alignment comparison.
Alignment Method PSNR $\uparrow$ MSE $\downarrow$ SSIM $\uparrow$ NCC $\uparrow$ AKD $\downarrow$
RealSR 13.87 4002.65 0.5072 0.3285 546.78
SIFT Alignment (Single time) 12.77 6332.79 0.4896 0.4313 546.78
[…]
@@
HAT 6670.32 1086.68 37.9%
DiffTSR 58502.14 8610.59 31.7%
DiT4SR w/o llava 160787.14 17385.14 23.7%
DreamClear w/o llava 412843.67 83193.81 21.4%
TADiSR 4497.96 342.32 36.6%
TIGER (ours) Stage 1 6215.85 663.5 43.0%